OpenStudy (anonymous):

Let f(x)= 1+ sin(x) a) Find the Linear Approximation of f(x) at the (c,f(c)) = (0,1) b) Evaluate both f(x) = 1+sin (x) and its linear approximation L(x) found in part(a) first as x = 0.01 and then at x=2. c) For which x value in part (b) is f(x) approximated with greater accuracy. I've tried to attempt the question but I'm not sure if I have the right answers. Please help!

OpenStudy (anonymous):

same as "find the equation for the line tangent to the graph"

OpenStudy (anonymous):

\[f(x)=1+\sin(x)\] \[f'(x)=\cos(x)\] \[m=f'(0)=\cos(0)=1\] \((x_1,y_1)\) is \((0,1)\) equation of line is \[y-1=1(x-0)\]or more simpley \[y=x+1\]

OpenStudy (anonymous):

Okay, that's what I got for the first one too. What about the other too? Do I have to just substitute?

OpenStudy (anonymous):

The other two*

OpenStudy (anonymous):

yes. evidently your approximation will be good at \(x=.01\) bur rather lousy at \(x=2\) since this is the linear approximation at \(x=0\)

OpenStudy (anonymous):

Yup, that's what I got as well. Thank you SO much! :)

OpenStudy (anonymous):

yw